# Splat the Net: Radiance Fields with Splattable Neural Primitives

### v1

@@ 6 Discussion and Conclusion
Although neural primitives exhibit substantial expressivity with limited memory resources, the complexity of the optimization landscape for millions of networks occasionally introduces convergence difficulties, hindering the expressivity of neural representations. A promising avenue for future research is to develop effective optimization or training strategies to fully unleash the expressivity of neural primitives. Moreover, as a general density field representation, our neural density field remains orthogonal to other techniques designed for color field and densification. Hence, integrating neural primitives with such advanced techniques is another interesting direction to explore in the future.
Appendix A hyperparameters
We train all scenes using the Adam optimizer (Diederik, 2014), with a learning rate of $10^{-3}$ for the MLP. For the datasets BlenderNeRF, Mip-NeRF360, Tanks & Temples, and Deep Blending, we adopt the following learning rate hyperparameters: primitive means ($1.6\times 10^{-4}$), scales ($5\times 10^{-3}$), quaternions ($10^{-3}$), and SH coefficients ($2.5\times 10^{-3}$). Population control is governed by a growing scale threshold of $10^{-2}$ and a pruning scale threshold of $0.5$.
Appendix B More Results
B.1 Numerical Results
BlenderNeRF Synthetic Dataset
Our neural primitive requires 41 parameters from its 8-neuron MLP, 10 from geometry (3 for means, 3 scales, and 4 for quaternion), and 48 from SHs, in total 99 parameters, $1.68\times$ more than 3DGS’ parameters (59). We report per-scene image metrics (PSNR, LPIPS, and SSIM) in Tab. 3, Tab. 4, and Tab. 5, under different memory budgets. For a fair comparison, we constrain the number of primitives in our system to half that of 3DGS and report numerical results under this setting. As shown in these three tables, the first two double rows exhibit an apple-to-apple comparison between 3DGS and our method under the same memory budget. Neural primitives outperform Gaussian primitives consistently, highlighting the expressivity of our representation. The last double rows in the three tables evaluate the performance of the two representations under unlimited memory budgets.
Real Datasets
Tables  6 and  7 present the per-scene performance metrics of our neural primitives, including PSNR, SSIM, LPIPS, the number of primitives, and the associated model memory footprint.
@@ B.1 Numerical Results
B.2 Visual Results
Primitive Expressivity
@@ Appendix C Applications
C.1 Volumetric Dynamic Novel View Synthesis
@@
The zero-order Spherical Harmonics (SH) coefficient of each primitive is modeled as a function of time $\xi_{t}$, expressed as a summation of a polynomial function and a Fourier series, similar to (Lin et al., 2024):
@@
Figure 11: More visual comparison on real datasets.
@@
We evaluate our method in the dynamic volumetric novel view synthesis setting using a synthetic dataset, including four volumetric effects from JangaFX***https://jangafx.com/software/embergen/download/free-vdb-animations. Each effect is recorded by 40 cameras on the upper hemisphere to capture temporal evolution, with 38 cameras for training and 2 for testing. The Colorful Smoke and Ground Explosion scenes contain 128 and 130 timesteps per camera, while Dust Tornado and Smoke Fire each have 100 timesteps per camera.
@@
$f_{\sigma}(\mathbf{x},\xi_{t})=W_{2}(\cos(W_{1}(\mathbf{x})+\xi_{t}\cdot W_{t}+\mathbf{b}_{1})+\mathbf{b}_{2}$ (13)
@@ Results
C.2 Relighting
@@
We evaluate our conduct relighting application using datasets provided by (Bi et al., 2024; Kang et al., 2019), including: (1) rendered images of synthetic NeRF scenes, and (2) rendered images of real captures.
@@ Results
Figure 12: We demonstrate relighting results using neural primitives.
Figure 13: We demonstrate our results for volumetric dynamic view synthesis. By introducing one additional dimension of time $\xi_{t}$, our neural primitives can reconstruct the scene’s evolution and synthesize coherent results for volumetric dynamic scenes.

### v2

@@ 6 Discussion and Conclusion
Although neural primitives exhibit substantial expressivity with limited memory resources, the complexity of the optimization landscape for millions of networks occasionally introduces convergence difficulties, hindering the expressivity of neural representations and slowing down convergence. For instance, on the MipNeRF360 dataset, neural primitives require an average wall-clock training time of 2.5 hours, which is approximately $2.5\times$ longer than 3DGS. A promising avenue for future research is to develop effective optimization or training strategies, such as the stochastic preconditioning technique  (Ling et al., 2025), to fully unleash the expressivity of neural primitives and accelerate convergence. Moreover, our neural density field is independent of the improvement strategies for the color field and densification in 3DGS, making it theoretically compatible with such techniques. For example, the error-guided adaptive densification strategies (Rota Bulò et al., 2024; Mallick et al., 2024) could be readily adapted to our framework. Spatially-varying textured primitives (Chao et al., 2025) could also be integrated with the proposed neural density field. Exploring such integrations remains an interesting future research direction.
Appendix A Integration Cost
Table 4: Comparison of GPU profiling metrics in Nsight Compute.
Metric 3DGS Ours
Compute (SM) Throughput [%] 84.52 80.02
Memory Throughput [%] 53.09 80.02
Executed Instructions [inst] 868,881,013 3,406,837,112
Goal
We conduct a comprehensive analysis of the computational overhead of neural and Gaussian primitives, from macro to micro-level evaluations on a single NVIDIA GeForce RTX 4090 GPU. We begin by evaluating the total rendering time under the same primitive budget, which provides the computational cost of the full rendering pipeline, including preprocessing, primitive sorting/tiling, integration, and alpha blending. We then analyze the computational cost in the render kernel function, which consists of the integration and blending processes. Finally, we focus on the integration step, decomposing it into the floating-point (FP) operation level.
Protocol
We perform computation analysis on a set of 100k primitives with random positions inside a volume and other random properties (MLP weights for neural primitives, opacities for gaussians) using our method and 3DGS, under the camera configurations of the synthetic NeRF dataset. We report FPS for total rendering time, similar to other 3DGS-based approaches. And we rely on Nsight Profiling to evaluate the render kernel function, providing compute and memory throughput, and executed instructions. We additionally manually decompose the integration process into floating-point (FP) operation level and report the FLOP number.
Analysis – FPS
Neural primitives achieve an FPS of 218.87, which is approximately $2.5\times$ slower than 3DGS (546.54). This gap is expected, as neural primitives require additional computation, including the MLP query, integration, and ray/ellipsoid intersection. In real scenes, our method needs $10\times$ fewer primitives, which balances computational complexity with the number of primitives, enabling real-time performance.
Analysis – Nsight
We report the Nsight profiling results by monitoring the render kernel function in Tab. 4, which consists of integration and alpha blending steps. Since our method and 3DGS employ the same alpha blending structure, the performance difference solely comes from the integration part. Both 3DGS and our method rely heavily on GPU compute ($84.52\%$ vs. $80.02\%$). Unlike 3DGS, which keeps its parameters in shared memory, our implementation must repeatedly access global memory for MLP weights, resulting in a higher memory throughput of $80.02\%$. The total number of executed instructions for neural primitives is $3.4\times 10^{9}$, approximately four times higher than that of 3DGS ($8.7\times 10^{8}$).
Figure 11: More visual comparison on real datasets.
@@ Appendix A Integration Cost
Analysis – FLOP
We further decompose the integration process and provide a FLOP-based analysis focused on the integration step. Using heuristic operation costs, where basic arithmetic, division, and exp/sin operations are assigned costs of 1, 4, and 30, respectively, 3DGS requires approximately 43 FLOPs per integration, whereas our method requires roughly 769 FLOPs.
Summary
Importantly, although neural primitives yield higher computational costs across FPS, Nsight, and FLOP-based analysis, our method requires roughly $10\times$ fewer primitives to represent the same scene as 3DGS. As a result, the overall rendering speed during inference remains over 100 FPS, which is comparable to 3DGS and ensures real-time performance.
Appendix B hyperparameters
We train all scenes using the Adam optimizer (Kingma and Ba, 2015), with a learning rate of $10^{-3}$ for the MLP. For the datasets BlenderNeRF, Mip-NeRF360, Tanks & Temples, and Deep Blending, we adopt the following learning rate hyperparameters: primitive means ($1.6\times 10^{-4}$), scales ($5\times 10^{-3}$), quaternions ($10^{-3}$), and SH coefficients ($2.5\times 10^{-3}$). Population control is governed by a growing scale threshold of $10^{-2}$ and a pruning scale threshold of $0.5$.
Appendix C More Results
C.1 Numerical Results
BlenderNeRF Synthetic Dataset
Our neural primitive requires 41 parameters from its 8-neuron MLP, 10 from geometry (3 for means, 3 scales, and 4 for quaternion), and 48 from SHs, in total 99 parameters, $1.68\times$ more than 3DGS parameters (59). We report per-scene image metrics (PSNR, LPIPS, and SSIM) in Tab. 5, Tab. 6, and Tab. 7, under different memory budgets. For a fair comparison, we constrain the number of primitives in our system to half that of 3DGS and report numerical results under this setting. As shown in these three tables, the first two double rows exhibit an apple-to-apple comparison between 3DGS and our method under the same memory budget. Neural primitives outperform Gaussian primitives consistently, highlighting the expressivity of our representation. The last double rows in the three tables evaluate the performance of the two representations under unlimited memory budgets.
Real Datasets
Tables  8 and  9 present the per-scene performance metrics of our neural primitives, including PSNR, SSIM, LPIPS, the number of primitives, and the associated model memory footprint.
C.2 Visual Results
Toy Examples
@@ Appendix D Applications
D.1 Volumetric Dynamic Novel View Synthesis
@@
The zero-order Spherical Harmonics (SH) coefficient of each primitive is modeled as a function of time $\xi_{t}$, expressed as a summation of a polynomial function and a Fourier series, similar to Lin et al. (2024):
@@
We evaluate our method in the dynamic volumetric novel view synthesis setting using a synthetic dataset, including four volumetric effects from JangaFX [^1] https://jangafx.com/software/embergen/download/free-vdb-animations. Each effect is recorded by 40 cameras on the upper hemisphere to capture temporal evolution, with 38 cameras for training and 2 for testing. The Colorful Smoke and Ground Explosion scenes contain 128 and 130 timesteps per camera, while Dust Tornado and Smoke Fire each have 100 timesteps per camera.
@@
$f_{\sigma}(\mathbf{x},\xi_{t})=W_{2}\,\cos(\omega_{0}\cdot W_{1}(\mathbf{x})+\xi_{t}\cdot W_{t}+\mathbf{b}_{1})+\mathbf{b}_{2}$ (13)
@@ Results
Figure 12: Relighting results under OLAT illumination using neural primitives.
Figure 13: We demonstrate our results for volumetric dynamic view synthesis. By introducing an additional time dimension $\xi_{t}$, our neural primitives can reconstruct the scene’s evolution and synthesize coherent results for volumetric dynamic scenes.
D.2 Relighting
@@
We conduct relighting application on One-Light-at-a-Time (OLAT) datasets provided by Bi et al. (2024); Kang et al. (2019), including: (1) rendered images of synthetic NeRF scenes, and (2) rendered images of real captures.